# DriveSuprim: Towards Precise Trajectory Selection for End-to-End Planning

### v2

@@ Abstract
In complex driving environments, autonomous vehicles must navigate safely. Relying on a single predicted path, as in regression-based approaches, usually does not explicitly assess the safety of the predicted trajectory. Selection-based methods address this by generating and scoring multiple trajectory candidates and predicting the safety score for each, but face optimization challenges in precisely selecting the best option from thousands of possibilities and distinguishing subtle but safety-critical differences, especially in rare or underrepresented scenarios. We propose DriveSuprim to overcome these challenges and advance the selection-based paradigm through a coarse-to-fine paradigm for progressive candidate filtering, a rotation-based augmentation method to improve robustness in out-of-distribution scenarios, and a self-distillation framework to stabilize training. DriveSuprim achieves state-of-the-art performance, reaching 93.5% PDMS in NAVSIM v1 and 87.1% EPDMS in NAVSIM v2 without extra data, demonstrating superior safety-critical capabilities, including collision avoidance and compliance with rules, while maintaining high trajectory quality in various driving scenarios.
1 Introduction
@@
End-to-end autonomous driving has traditionally relied on regression-based approaches that predict a single trajectory to mimic expert behavior prakash2021multi ; chitta2021neat ; chen2021learning ; jiang2023vad ; hu2023planning ; diffusiondrive ; hu2023gaia ; wang2023drivedreamer . While regression is a common approach, it fundamentally lacks the ability to evaluate multiple alternatives in safety-critical scenarios where subtle trajectory differences can significantly impact outcomes.
@@
In recent years, selection-based methods chen2024vadv2 ; li2024hydra ; li2024hydramdp_pp ; wang2025enhancing have clearly outperformed regression approaches. The key is their capability to generate and explicitly evaluate diverse trajectory candidates using comprehensive safety metrics such as collision risk and rule compliance dauner2024navsim . This explicit comparison enables the system to select the safest and most appropriate trajectory from multiple alternatives, addressing safety-critical issues that regression-based methods cannot address.
Our oracle study demonstrates the substantial potential of selection-based methods: when making optimal selection, these approaches can even surpass human demonstrations in safety-critical metrics (see Tab 1). This performance ceiling highlights why selection-based planning has become the preferred paradigm for autonomous driving systems requiring robust safety guarantees.
Figure 1: The overall pipeline of our method. Selection-based methods struggle to distinguish suboptimal “hard negative” trajectories, perform poorly in turning, and utilize hard binary labels for scoring training. Our proposed DriveSuprim introduces a coarse-to-fine refinement framework and a rotation-based data augmentation method with self-distillation to address these weaknesses. The green trajectory in the image is the ground-truth trajectory, the red and orange trajectories are obviously unsafe and seemingly correct trajectories, and the blue trajectory denotes the refined trajectory predicted by our model.
However, selection-based methods still face three critical limitations: First, selection-based methods struggle to differentiate between optimal trajectories and similar but suboptimal alternatives. This weakness makes the ideal perfect trajectory selection in Tab 1 challenging to implement and limits the model performance. During training, these models encounter thousands of trajectory candidates where the vast majority are unsafe or impractical (“easy negatives”, the red trajectory in Fig 1). These easy-to-reject options dominate the training process, causing the model to primarily learn to avoid clearly bad choices. As a result, the model receives inadequate training signals for distinguishing between reasonable-looking trajectories with subtle but important differences (“hard negatives”, the orange trajectory in Fig 1). The overwhelming number of obvious negative examples hinders the model’s ability to develop fine-grained discrimination capability, which is crucial for selecting optimal trajectories when presented with multiple plausible options. Such discrimination capability is essential for safely navigating complex driving scenarios.
Second, selection-based methods suffer from directional bias in trajectory distribution. This bias manifests as an imbalance in training data that, while reflecting real-world driving patterns where straight driving predominates, leads to models that perform relatively poorly in turning scenarios. Training with such imbalanced data naturally results in models that excel at straight-line driving but struggle with turns and complex maneuvers. Even advanced autonomous driving datasets like NAVSIM dauner2024navsim exhibit this limitation. Our analysis reveals that only 8% of ground-truth trajectories in NAVSIM involve turns exceeding 30 degrees. While this distribution may reflect typical driving patterns, it creates a significant challenge for learning models, which require sufficient examples of all maneuver types to develop robust capabilities. This directional bias significantly impairs the model’s ability to learn from and correctly execute large-angle turning trajectories, particularly in navigation-critical scenarios where turns are essential.
Third, selection-based methods typically rely on binary classification for safety-related decisions, labeling trajectories as "safe" or "unsafe" based on specific thresholds for collision risk or rule
@@
compliance. This binary approach creates hard decision boundaries where trajectories just above or below a safety threshold could be treated entirely differently. As a result, models become overly sensitive to minor changes in trajectory features, which causes inconsistent behavior when slight variations in collision risk or rule adherence could suddenly thoroughly flip a trajectory from being selected to rejected.
We present DriveSuprim, a novel method that tackles these three critical challenges in selection-based trajectory prediction and makes more precise trajectory selection. Our solution replaces rigid binary classifications with probabilistic soft-label distillation, enabling nuanced decision-making in complex driving scenarios. Our contributions include:
@@
  • We propose an integrated pipeline combining rotation-based data augmentation with self-distillation to address directional bias and hard decision boundaries. Our approach synthesizes challenging turning scenarios while leveraging teacher-generated soft pseudo-labels, effectively balancing trajectory distributions as shown in Fig 4 and realizing better optimization for training a more robust model.
@@
  • DriveSuprim achieves state-of-the-art performance on the NAVSIM benchmark purely based on the public dataset, demonstrating the effectiveness of our model in handling challenging driving scenarios. Our approach achieves state-of-the-art results on NAVSIMv1 and NAVSIMv2, significantly outperforming the previous methods by 3.6% and 1.5%.
2 Related Works
2.1 End-to-end planning
Autonomous driving has traditionally relied on modular pipelines that separate perception from planning. However, UniAD hu2023planning highlights several limitations of this approach, including information loss and error propagation. To address these challenges, end-to-end driving methods chen2020learning ; chitta2021neat ; prakash2021multi ; chen2021learning ; wu2022trajectory ; shao2023safety ; shao2023reasonnet ; chitta2022transfuser ; jia2023think ; jaeger2023hidden ; zhang2023coaching ; hu2022st ; hu2023planning ; jiang2023vad ; li2024ego ; weng2024para ; ye2023fusionad ; chen2024vadv2 ; li2024hydra ; toromanoff2020end ; zhang2021end ; wang2025enhancing unify the perception-to-planning pipeline within a single optimizable network. Many of these models process raw sensor inputs and directly output driving trajectories. While some methods chen2021learning ; toromanoff2020end ; zhang2021end use reinforcement learning (RL) to learn through interaction with simulated environments, the majority adopt imitation learning (IL), training from expert demonstrations without environment interaction. Most IL-based approaches chitta2022transfuser ; hu2023planning ; jiang2023vad ; li2025finetuning ; liao2024diffusiondrive generate a single trajectory using regression or diffusion-based methods to mimic expert behavior.
More recently, selection-based methods jiang2023vad ; chen2024vadv2 ; li2024hydra ; li2024hydramdp_pp ; wang2025enhancing have emerged. These models evaluate a diverse set of candidate trajectories by scoring them against safety-focused metrics (e.g., PDM scores dauner2024navsim ). A prominent example is Hydra-MDP li2024hydra , which won the recent NAVSIM challenge. It employs multiple rule-based teachers and distills them into the planner to create diverse trajectory candidates tailored to different evaluation metrics.
Our proposed model also falls under the selection-based paradigm. However, unlike prior methods that perform a single-shot selection from a fixed candidate set—often leading to suboptimal decisions—we introduce a coarse-to-fine selection and refinement strategy. This approach significantly improves selection precision by progressively narrowing down the trajectory set to the most optimal candidates.
2.2 Iterative & Multi-stage Refinement
Iterative refinement has been widely adopted to improve results in optical flow ilg2017flownet ; ranjan2017optical ; sun2019models ; yang2019volumetric ; hui2018liteflownet ; teed2020raft ; xu2022gmflow , motion estimation sun2024refining ; harley2022particle ; zheng2023pointodyssey , and related tasks liang2018learning ; zhou2018deeptam ; teed2018deepv2d ; li2018recurrent . A common strategy in these works is to iteratively propagate features or trajectory estimates through a shared module to progressively refine the predictions. Inspired by this, iterative refinement has also been applied in object detection zhu2020deformable ; cai2018cascade to improve detection performance. For instance, in Deformable DETR, each decoder layer refines bounding boxes based on predictions from the previous layer.
In our work, we similarly adopt a multi-stage refinement strategy to improve trajectory selection accuracy. However, a key distinction lies in the mechanism of refinement: rather than repeatedly updating fixed-dimensional features, we implement selection from a fixed vocabulary of candidate trajectories. After selection, the search space is progressively narrowed to a more precise subset, improving the final prediction quality.
2.3 Augmentation for Enhanced Robustness
Robustness has long been a critical focus in computer vision research croce2020robustbench ; ganin2016domain ; sun2022benchmarking ; tang2021robustart ; yuan2023robust ; rebuffi2021data . Early studies demonstrated that image models are highly sensitive to minor domain shifts azulay2019deep , adversarial perturbations szegedy2013intriguing ; goodfellow2014explaining , and common real-world corruptions such as brightness variations and fog pei2017towards ; hendrycks2018benchmarking . For example, MNIST-C mu2019mnist introduced 15 distinct corruption types to benchmark model performance against diverse failure modes. Motivated by these insights, several methods rusak2020simple ; mintun2021interaction ; saikia2021improving ; kar20223d utilize corruption-based augmentations, such as adding Gaussian and speckle noise, to enhance robustness. Inspired by this established research in image robustness, our study is the first to explore the use of similar corruption-based augmentation techniques specifically for end-to-end driving models. We introduce targeted perturbations and corruptions tailored to autonomous driving scenarios, addressing critical domain shifts—particularly the overrepresentation of straightforward driving trajectories—which pose challenges for scenarios involving complex maneuvers such as turns.
3 Methods
We introduce our method in this section. Firstly, we introduce some preliminaries, including the end-to-end planning and the selection-based planning method. Next, we introduce our proposed coarse-to-fine selection paradigm, rotation-based data augmentation method and self-distillation.
3.1 Preliminaries
@@
In the autonomous driving domain, the end-to-end planning requires the planning system to output a future trajectory $T$ based on input sensor data, like RGB image or Lidar point cloud:
@@
Selection-based planning
The selection-based planning paradigm predefines a trajectory vocabulary $\{\tau_{i}\}_{i=1}^{N}$, which covers $N$ planning trajectories. Given a specific driving scenario, the quality of each trajectory can be measured by several metrics, like the $l_{2}$ distance to the human teacher trajectory, or metrics considering driving safety and traffic rule adherence. The selection-based planning paradigm learns a scorer that generates trajectory scores $\{s_{i}\}_{i=1}^{N}$ revealing trajectory quality. The trajectory with the highest score is chosen as the prediction result in inference:
$T=\tau_{k},\quad\text{where }k=\arg\max_{i}s_{i}$ (2)
3.2 Coarse-to-Fine Trajectory selection
DriveSuprim proposes a coarse-to-fine trajectory selection paradigm comprising coarse filtering and fine-grained scoring, improving model capability on distinguishing hard negative trajectories, as shown in Fig 2. In the coarse filtering stage, the model selects a series of trajectory candidates based on the predicted scores, similar to classic selection-based approaches. The fine-grained scoring stage then produces more accurate scores for the filtered trajectories.
@@ 3.2 Coarse-to-Fine Trajectory selection
The coarse filtering stage scores all trajectories in the vocabulary and filters a smaller set of candidates for the next stage. Here we apply the same strategy as the previous selection-based method li2024hydra . Each trajectory is encoded into a vector by a lightweight MLP encoder, and then the trajectory feature cross-attends with the input image feature to extract the planning-related information. To measure the quality of each trajectory in a specific scenario, several prediction heads are applied on the refined trajectory feature to regress the $l_{2}$ distance to the ground-truth human trajectory and the rule-based metric scores:
$\mathcal{E}_{\mathrm{img}}=\mathrm{Enc_{i}}(I),\quad f_{j}=\mathrm{Enc_{t}}(\tau_{j}),\quad g_{j}=\mathrm{TransDec}(\mathcal{E}_{\mathrm{img}},f_{j})$ (3)
@@
where $\mathrm{Enc_{i}}$ and $\mathrm{Enc_{t}}$ are image encoder and trajectory encoder, $I$ and $\mathcal{E}_{\mathrm{img}}$ are the input image and image feature, $\tau_{j}$ and $f_{j}$ denote the trajectory and the encoded trajectory feature, $\mathrm{TransDec}$ denotes the Trajectory Decoder, which is a Transformer decoder vaswani2017attention , $g_{j}$ denotes the refined trajectory feature, $\mathrm{head^{(m)}}$ denotes the prediction head of metric $m$, and $s_{j}^{(m)}$ denotes the score of trajectory $\tau_{j}$ on the evalutation metric $m$.
@@
In this stage, a Transformer decoder similar to the first stage is applied to make fine-grained scoring and further distinguish the trajectories in the last-stage filtered candidates.
Specifically, we obtain the refined score from the $l$-th decoder layer output feature, and optimize with the ground truth trajectory score:
@@
where $\mathrm{RefineDec}$ is a Transformer decoder with $n_{\mathrm{ref}}$ layers, $h_{j,l}$ is the refined feature output from the $l$-th layer of the Refinement Transformer, $s_{j,l}^{(m)}$ denotes the score of $\tau_{j}$ on metric $m$ output by the $l$-th decoder layer. The output trajectory $T$ is selected based on the last-layer output score $s_{j,n_{\mathrm{ref}}}^{(m)}$. This method can ease the difficulties of selecting the best trajectory from the top-k trajectories containing a large proportion of hard negative candidates.
The loss for our coarse-to-fine module on the original dataset is represented as:
@@
where $L_{\mathrm{coarse}}$ and $L_{\mathrm{refine}}$ respectively train the coarse filtering stage and the fine-grained scoring stage. $L_{\mathrm{coarse}}$ consists of an imitation loss $L_{\mathrm{imi}}$ li2024hydra ; li2024hydramdp_pp and a binary cross-entropy between the predicted trajectory score and the ground truth metric score:
@@
where $y_{i}^{(m)}$ is the ground-truth metric score of each trajectory $\tau_{i}$ on metric $m$, and $s_{i}^{(m)}$ is its corresponding predicted score. The fine-grained scoring stage loss $L_{\mathrm{refine}}$ is similar to $L_{\mathrm{coarse}}$, while it only considers the filtered trajectories $T_{\mathrm{filter}}$ rather than the entire vocabulary, and is applied on each decoder layer output.
3.3 Rotation-based data augmentation
@@
To mitigate the data imbalance, our model introduces an end-to-end rotation-based augmentation pipeline, where a 2D horizontal view transformation is applied to the sensor input data to simulate the ego-vehicle rotation in the 3D space. This approach conveniently synthesizes more challenging scenarios and diversifies the driving scenarios, enabling the model to learn to output accurate trajectories regardless of the vehicle orientation.
@@
As shown in Fig 3, for each scenario, we sample a random angle $\theta$ from a uniform distribution $U[-\Theta,\Theta]$, where $\Theta$ is the angle boundary. The positive angle indicates the leftward rotation of the ego vehicle. To prepare the input RGB image, camera images corresponding to the original field of view (FOV) along with images from two extended views are concatenated to simulate a “pseudo panoramic view”, then the input image is cropped from the concatenated image according to a shifting window based on $\theta$. For example, for a 1-camera FOV input setting, we choose these three cameras as our input: $f$ (front camera), $l_{0}$ (front-left camera), and $r_{0}$ (front-right camera), where $f$ corresponds to the original field of view, and $l_{0}$ and $r_{0}$ correspond to the extended view.
@@
The ground truth human trajectory $(u_{1},u_{2},\dots,u_{l})$ in the rotated scenario is generated by a trivial rotation approach. Under a specific rotation angle $\theta$, each location $u_{j}$ of the human trajectory $T_{h}$ applies a 2D-rotation transformation surrounding the initial vehicle position $u_{0}$, and the rotation angle is $-\theta$. Given augmented camera views and the prediction of our model, we calculate a loss $L_{\mathrm{aug}}$ for training, which has the same formulation as $L_{\mathrm{ori}}$.
3.4 Self-distillation with Soft-labeling
@@
Instead of training the selection-based model to fit a hard decision boundary given limited environmental contexts (e.g., insufficient fields of view), which can impair training stability, we propose a self-distillation framework with teacher-generated soft labels to stabilize training. The self-distillation framework consists of a teacher and a student model, both sharing the identical architecture. The student parameters are updated via standard gradient descent, whereas the teacher parameters are updated using an exponential moving average (EMA) of the student’s parameters.
@@
During training, the student receives both original and augmented data to calculate $L_{\mathrm{ori}}$ and $L_{\mathrm{aug}}$, as shown in Fig 2. The teacher only receives original data to generate scores served as soft labels, where a threshold $\delta_{m}$ is introduced to control the difference between the teacher output and the ground-truth:
@@
where $s_{i,\mathrm{teacher}}^{(m)}$ is the score of trajectory $i$ on metric $m$ predicted by the teacher, $y_{i}^{(m)}$ is the ground-truth score of trajectory $i$ on metric $m$, $\mathrm{clip}\left(\cdot,\alpha,\beta\right)$ denotes clipping the input value to the interval $[\alpha,\beta]$, and $\hat{y}_{i}^{(m)}$ is the soft label, and $L_{\mathrm{imi-soft}}$ is a soft version of $L_{\mathrm{imi}}$. We shift the human trajectory toward the teacher model’s output trajectory for at most 1 meter, and adopt this new trajectory to calculate the imitation loss $L_{\mathrm{imi-soft}}$. The overall training loss of the student model is represented as:

### v3

@@ Abstract
Autonomous vehicles must navigate safely in complex driving environments. Imitating a single expert trajectory, as in regression-based approaches, usually does not explicitly assess the safety of the predicted trajectory. Selection-based methods address this by generating and scoring multiple trajectory candidates and predicting the safety score for each. However, they face optimization challenges in precisely selecting the best option from thousands of candidates and distinguishing subtle but safety-critical differences, especially in rare and challenging scenarios. We propose DriveSuprim to overcome these challenges and advance the selection-based paradigm through a coarse-to-fine paradigm for progressive candidate filtering, a rotation-based augmentation method to improve robustness in out-of-distribution scenarios, and a self-distillation framework to stabilize training. DriveSuprim achieves state-of-the-art performance, reaching 93.5% PDMS in NAVSIM v1 and 87.1% EPDMS in NAVSIM v2 without extra data, with 83.02 Driving Score and 60.00 Success Rate on the Bench2Drive benchmark, demonstrating superior planning capabilities in various driving scenarios.
Codehttps://github.com/William-Yao-2000/DriveSuprim
Introduction
@@
End-to-end autonomous driving has traditionally relied on regression-based approaches that predict a single trajectory to mimic expert behavior  (Jiang et al. 2023; Hu et al. 2023; Wang et al. 2023) . While regression is a common approach, it fundamentally lacks the ability to evaluate multiple alternatives in safety-critical scenarios where subtle trajectory differences can significantly impact outcomes.
@@
In recent years, selection-based methods (Chen et al. 2024; Li et al. 2024c, a; Wang et al. 2025) have clearly outperformed regression approaches. The key is their capability to generate and evaluate diverse trajectory candidates using comprehensive safety metrics such as collision risk and driving rule compliance (Dauner et al. 2024). This explicit comparison enables the system to select the safest and most appropriate trajectory from multiple alternatives, addressing safety-critical issues that regression-based methods cannot address. Our oracle study in Tab 1 demonstrates the substantial potential of selection-based methods: when making ideal optimal selection, these approaches can even surpass human demonstrations in NAVSIM (Dauner et al. 2024) safety-critical metrics. This performance ceiling highlights why selection-based planning has become the preferred paradigm for autonomous driving systems requiring robust safety guarantees.
@@
Figure 1: Overall pipeline of our method. Selection-based methods struggle to distinguish suboptimal trajectories, perform poorly in turning, and utilize hard binary labels in training. DriveSuprim introduces coarse-to-fine refinement and a rotation-based data augmentation method with self-distillation to address these weaknesses. The green trajectory is the ground-truth trajectory, and the red and orange trajectories are obviously unsafe and seemingly correct trajectory candidates in the trajectory vocabulary.
However, selection-based methods still face three critical limitations, which make them difficult to reach the ideal perfect trajectory selection displayed in Tab 1 and limit the model performance. First, selection-based methods struggle to distinguish the optimal trajectory from similar but suboptimal alternatives. During training, the model encounters thousands of trajectory candidates where the vast majority are unsafe or impractical (“easy negatives”, the red trajectory in Fig 1). These easy-to-reject options dominate the training process and gradient, causing the model to focus primarily on avoiding obviously incorrect choices. In contrast, the model receives insufficient supervision to select the most suitable trajectory from reasonable-looking trajectories with subtle but important differences (“hard negatives”, the orange trajectory in Fig 1). The overwhelming number of obvious negative examples hinders the model’s ability to develop fine-grained discrimination capability, which is crucial for selecting optimal trajectories when presented with multiple plausible options, to avoid route deviation or collision.
Second, selection-based methods suffer from directional bias in trajectory distribution. This bias manifests as an imbalance in training data that, while reflecting real-world driving patterns where straight driving predominates, leads to models that perform relatively poorly in turning scenarios. Training with such imbalanced data naturally results in models that excel at straight-line driving but struggle with turns and complex maneuvers. Even advanced autonomous driving datasets like NAVSIM exhibit this limitation. We find that only 18% of ground-truth trajectories in NAVSIM involve turns exceeding 30 degrees. While this distribution may reflect typical driving patterns, it creates a significant challenge for learning models, which require sufficient examples of all maneuver types to develop robust capabilities. This directional bias significantly impairs the model’s ability to select the most precise large-angle turning trajectories, particularly in navigation-critical scenarios where turns are essential.
Third, selection-based methods typically rely on binary classification for safety-related decisions, labeling trajectories as “safe” or “unsafe” based on specific thresholds for collision risk or rule compliance. This binary approach creates hard decision boundaries where trajectories just above or below a safety threshold could be treated entirely differently. Slight variations in score could suddenly thoroughly flip a trajectory label from being selected to being rejected. As a result, models risk becoming overly sensitive to minor changes in trajectory features, which causes inconsistent behavior.
We present DriveSuprim, a novel method that tackles these three critical challenges in selection-based trajectory prediction and makes more nuanced and precise trajectory selection. Our contributions include:
@@
  • We propose an integrated training pipeline combining rotation-based data augmentation with self-distillation to address directional bias and hard decision boundaries. Our approach synthesizes challenging turning scenarios and leverages teacher-generated soft pseudo-labels, effectively balancing trajectory distributions and realizing better optimization for training a more robust model.
@@
  • DriveSuprim achieves state-of-the-art performance on the NAVSIM and Bench2Drive benchmark, demonstrating the effectiveness of our model in handling challenging driving scenarios. DriveSuprim significantly outperforms the previous methods by 3.6% and 1.5% on NAVSIMv1 and NAVSIMv2 without introducing other training data. On the Bench2Drive benchmark, our method achieves 83.02 and 60.00 on driving score and success rate.
Related Works
@@
Autonomous driving has traditionally relied on modular pipelines that separate perception from planning. However, UniAD (Hu et al. 2023) highlights several limitations of this approach, including information loss and error propagation. To address these challenges, end-to-end driving methods  (Chen et al. 2019; Chen, Koltun, and Krähenbühl 2021; Chitta et al. 2022; Hu et al. 2023; Jiang et al. 2023; Li et al. 2024d, c; Wang et al. 2025) unify the perception-to-planning pipeline within a single optimizable network. They process raw sensor inputs and directly output driving trajectories. While some methods  (Chen, Koltun, and Krähenbühl 2021; Zhang et al. 2021) use reinforcement learning (RL) to learn through interaction with simulated environments, the majority adopt imitation learning (IL), training from expert demonstrations without interaction. Most IL-based approaches  (Chitta et al. 2022; Hu et al. 2023; Jiang et al. 2023; Li et al. 2025a; Liao et al. 2025) generate a single trajectory using regression or diffusion-based methods to mimic expert behavior.
More recently, selection-based methods (Chen et al. 2024; Li et al. 2024c, a; Wang et al. 2025) have emerged. These models evaluate a diverse set of candidate trajectories by scoring them on safety-focused metrics (e.g., PDM scores (Dauner et al. 2024)). A prominent example is Hydra-MDP (Li et al. 2024c), which employs multiple rule-based teachers and distills them into the planner to create diverse trajectory candidates tailored to different evaluation metrics.
Our proposed model also falls under the selection-based paradigm. However, unlike prior methods that perform a single-shot selection from a fixed candidate set—probably leading to suboptimal decisions—we introduce a coarse-to-fine selection and refinement strategy. This approach significantly improves selection precision by progressively narrowing down the trajectory set to the most optimal candidates.
Iterative & Multi-stage Refinement
Iterative refinement has been widely adopted to improve results in optical flow  (Ilg et al. 2017; Hui, Tang, and Loy 2018; Teed and Deng 2020; Xu et al. 2022) and motion estimation  (Sun, Harley, and Guibas 2024; Zheng et al. 2023). A common strategy in these works is to iteratively propagate features or trajectory estimates through a shared module to progressively refine the predictions. Inspired by this, iterative refinement is also applied in object detection (Zhu et al. 2020; Cai and Vasconcelos 2018) to improve performance. For instance, in Deformable DETR, each decoder layer refines bounding boxes based on predictions from the previous layer.
We similarly adopt a multi-stage refinement strategy to improve trajectory selection accuracy. However, rather than repeatedly updating fixed-dimensional features, we implement selection from a fixed vocabulary of candidate trajectories. After selection, the search space is progressively narrowed to a more precise subset, improving the final prediction quality.
Augmentation for Enhanced Robustness
Robustness has long been a critical focus in computer vision research  (Ganin et al. 2016; Croce et al. 2020; Rebuffi et al. 2021; Sun et al. 2022). Early studies demonstrated that image models are highly sensitive to minor domain shifts (Azulay and Weiss 2019) and adversarial perturbations (Szegedy 2013; Goodfellow, Shlens, and Szegedy 2014). For example, MNIST-C (Mu and Gilmer 2019) introduces 15 distinct corruption types to benchmark model performance against diverse failure modes. Motivated by these insights, several methods  (Rusak et al. 2020; Mintun, Kirillov, and Xie 2021; Kar et al. 2022) utilize corruption-based augmentations, such as adding Gaussian and speckle noise, to enhance robustness. Inspired by these methods, our study explores the use of similar corruption-based augmentation techniques specifically for end-to-end driving models. We introduce targeted perturbations tailored to autonomous driving scenarios, addressing critical domain shifts—particularly the overrepresentation of straightforward driving trajectories—which pose challenges for scenarios involving complex maneuvers such as turns.
Methods
We introduce our method in this section. Firstly, we introduce preliminaries about the end-to-end planning and the selection-based method. Next, we introduce our proposed coarse-to-fine selection paradigm, rotation-based data augmentation method and self-distillation framework.
Preliminaries
End-to-End planning & Selection-based planning
In autonomous driving, the end-to-end planning requires the planning system to output a future trajectory $T$ based on input sensor data, like RGB image or Lidar point cloud:
@@ Methods
Among the end-to-end planning methods, the selection-based paradigm predefines a trajectory vocabulary $\{\tau_{i}\}_{i=1}^{N}$ covering $N$ planning trajectories. Given a specific driving scenario, the quality of each trajectory is evaluated by several metrics, like the $l_{2}$ distance to the human teacher trajectory, or metrics considering driving safety and traffic rule adherence. The model learns a scorer that generates trajectory scores $\{s_{i}\}_{i=1}^{N}$ revealing trajectory quality. The trajectory $\tau_{k}$ with the highest score is chosen as the predicted result in inference.
Coarse-to-Fine Trajectory selection
DriveSuprim proposes a coarse-to-fine trajectory selection paradigm comprising coarse filtering and fine-grained scoring, improving model capability in distinguishing hard negative trajectories. As shown in Fig 2, in the coarse filtering stage, the model selects several trajectory candidates based on predicted scores, similar to classic selection-based approaches. The fine-grained scoring stage then produces more accurate scores for the filtered trajectories.
@@
The coarse filtering stage scores all trajectories in the vocabulary and filters a smaller set of candidates for the next stage. Here we apply the same strategy as the previous selection-based method (Li et al. 2024c). The trajectory feature cross-attends with the image feature to extract planning-related information, then several prediction heads are applied on the refined trajectory feature to regress the normalized $l_{2}$ distance to the ground-truth human trajectory and the rule-based metric scores:
$\mathcal{E}_{\mathrm{img}}=\mathrm{Enc_{i}}(I),f_{j}=\mathrm{Enc_{t}}(\tau_{j}),$ (2)
$g_{j}=\mathrm{TransDec}(\mathcal{E}_{\mathrm{img}},f_{j})$ (3)
@@
where $\mathrm{Enc_{i}}$ and $\mathrm{Enc_{t}}$ are image encoder and trajectory encoder, $I$ and $\mathcal{E}_{\mathrm{img}}$ are the input image and image feature, $\tau_{j}$ and $f_{j}$ denote the trajectory and the encoded trajectory feature, $\mathrm{TransDec}$ denotes the Trajectory Decoder, which is a Transformer decoder (Vaswani et al. 2017), $g_{j}$ denotes the refined trajectory feature, $\mathrm{head^{(m)}}$ denotes the prediction head of evalutation metric $m$, $s_{j}^{(m)}$ denotes the prediction score of trajectory $\tau_{j}$ on metric $m$.
@@
In this stage, a Transformer decoder similar to the first stage is applied to make fine-grained scoring and further distinguish the trajectories in the first-stage filtered candidates, which contain a large proportion of hard negatives. Specifically, we obtain the refined score from the $l$-th decoder layer output feature, and optimize with the ground truth trajectory score:
@@
where $\mathrm{RefineDec}$ is a Transformer decoder with $n_{\mathrm{ref}}$ layers, $h_{j,l}$ is the refined feature of $\tau_{j}$ from the $l$-th layer of the Refinement Transformer, $s_{j,l}^{(m)}$ denotes the score of $\tau_{j}$ on metric $m$ output by the $l$-th decoder layer. The output trajectory $\tau_{k}$ is selected based on the highest last-layer output score $s_{k,n_{\mathrm{ref}}}^{(m)}$.
The loss for the coarse-to-fine module on the original dataset is represented as:
@@
where $L_{\mathrm{coarse}}$ and $L_{\mathrm{refine}}$ respectively train the coarse filtering stage and the fine-grained scoring stage. $L_{\mathrm{coarse}}$ consists of an imitation loss $L_{\mathrm{imi}}$ (Li et al. 2024c, a) and a binary cross-entropy between the predicted trajectory score and the ground truth metric score:
@@
where $y_{i}^{(m)}$ and $s_{i}^{(m)}$ are the ground-truth metric score and the predicted score of trajectory $\tau_{i}$ on metric $m$. $L_{\mathrm{refine}}$ is similar to $L_{\mathrm{coarse}}$, while it only considers the filtered trajectories $T_{\mathrm{filter}}$ rather than the entire vocabulary, and is applied on each decoder layer output.
Rotation-based data augmentation
@@
To mitigate the data imbalance, we introduce an end-to-end rotation-based augmentation pipeline, where a 2D horizontal view transformation is applied to the sensor input data to simulate the ego-vehicle rotation in the 3D space. This approach easily synthesizes more challenging scenarios and diversifies the driving scenarios, enabling the model to precisely select trajectories regardless of the vehicle orientation.
@@
As shown in Fig 3, for each scenario, we sample a random angle $\theta$ from a uniform distribution $U[-\Theta,\Theta]$, where $\Theta$ is the angle boundary. The positive angle indicates the leftward rotation of the ego vehicle. Camera images corresponding to the original field of view (FOV) along with images from two extended views are concatenated to simulate a “pseudo panoramic view”, then the input image is cropped from the concatenated image according to a shifting window based on $\theta$. For example, for a 1-camera FOV input setting, we choose these three cameras as our input: $f$ (front), $l_{0}$ (front-left), and $r_{0}$ (front-right), where $f$ corresponds to the original field of view, and $l_{0}$ and $r_{0}$ correspond to the extended view.
@@
In the synthesized rotated scenario, the ground truth human trajectory $(u_{1},u_{2},\dots,u_{l})$ is generated by a trivial rotation approach: under a specific rotation angle $\theta$, each location $u_{j}$ of the human trajectory $T_{h}$ applies a 2D-rotation transformation surrounding the initial vehicle position $u_{0}$, and the rotation angle is $\theta$, ensuring the world coordinate of the trajectory remains unchanged. Given augmented camera views and the prediction of our model, we calculate a loss $L_{\mathrm{aug}}$ for training, which has the same formulation as $L_{\mathrm{ori}}$.
Self-distillation with Soft-labeling
@@
Instead of training the selection-based model to fit a hard decision boundary of trajectory evaluation, which can impair training stability, we propose a self-distillation framework with teacher-generated soft labels to stabilize model training. The self-distillation framework consists of a teacher and a student model, both sharing the same architecture. The student is updated via standard gradient descent, whereas the teacher is updated using an exponential moving average (EMA) of the student’s parameters.
@@
During training, the student receives noisier input, including both original and augmented data to calculate $L_{\mathrm{ori}}$ and $L_{\mathrm{aug}}$, as shown in Fig 2. The teacher only receives original data to generate scores served as soft labels, where a clipping threshold $\delta_{m}$ is introduced to control the gap between the teacher output and the ground-truth:
@@
where $s_{i,\mathrm{teacher}}^{(m)}$ is the score of trajectory $i$ on metric $m$ predicted by the teacher, $y_{i}^{(m)}$ is the ground-truth score of trajectory $i$ on metric $m$, $\mathrm{clip}\left(\cdot,\alpha,\beta\right)$ denotes clipping the input value to the interval $\alpha,\beta]$, and $\hat{y}_{i}^{(m)}$ is the soft label. $L_{\mathrm{imi-soft}}$ is a soft version of $L_{\mathrm{imi}}$, where the human trajectory is shifted toward the teacher model’s output trajectory for at most 1 meter, and is adopted to calculate the imitation loss. The overall training loss of the student model is: